OpenStudy (anonymous):

can someone help me find the derivative of 4xy+(−4)y+5x=6 by implicit differentiation.

OpenStudy (anonymous):

let us do this! Where would you try and start?

OpenStudy (mathmale):

The +(-4)y in 4xy+(−4)y+5x=6 is a bit unusual. Are you sure it's not just 4xy−4y+5x=6 ??

OpenStudy (anonymous):

It's probably -4y instead of +(-4)y

OpenStudy (mathmale):

Can you go back to the original problem and double-check, before we invest time in finding the derivative of this implicit function?

OpenStudy (anonymous):

It is +(-4)y but that's odd

OpenStudy (anonymous):

so... 4xy−4y+5x=6. you are able to take the derivative of 4xy yes? and also the rest of this equation?

OpenStudy (anonymous):

I figured it out!! Thank you guys